1) Find the wavelength of a sound wave which has a frequency of 410 Hz. The speed of sound is 343 m/s.

2) Consider light with a wavelength equal to 576 nm. The speed of light is 3.00
Answer #1

1,

f = 410 Hz

v= 343 m/s

Now,

wavelength = v/f = 343 / 410 = 0.8365 m

2.

wavelength = 576 nm = 576 x 10^-9 m

v = 3 x 10^8 m/s

(a) wavelength = v/f

f = v / wavelength

f = 3 x 10^8 / 576 x 10^-9

f = 5.2 x 10^14 Hz

Time Period (T) = 1 / f = 1/(5.2 x 10^14) = 1.92 x 10^-15 sec = 1.92 femtosecond.

3.

(a)

wavelength = v/f = 3 x 10^8 / 935 x 10^3 = 320.85 m

(b)

wavelength = v / f = 3 x 10^8 / 87.2 x 10^6 = 3.44 m
Answer #3

1) wl = v/f = 343/410 = 0.8365 m

2) a) f = 3x10^8 / 576x10^-9 = 5.2x10^14 Hz

b) 1/f = 1/(5.2x10^14) = 1.92x10^-15 sec

3) a) Frequency f= 935 kHz = 935 * 10 ^ 3 Hz

speed of light c = 3 * 10 ^ 8 m / s
wavelength w = c / f = 320.86 m
b) frequency f ' = 87.2 * 10 ^ 6 Hz
wavelength w' = c / f ' = 3.44 m
Answer #4

QUESTION 1) HERE WE NEED TO FIND THE VALUE OF THE WAVELENGTH .

LAMBDA = v/f = 343 m/s / 410 hz = 0.836585365 m ===================PART A) ANSWER)

QUESTION 2) INCOMPLETE INFORMATION PROVIDED.